OpenStudy (anonymous):

What is the center and radius of the circle with the equation (x + 10)2 + (y + 15)2 = 121 ?

jimthompson5910 (jim_thompson5910):

Hint: (x + 10)^2 + (y + 15)^2 = 121 (x - (-10))^2 + (y - (-15))^2 = 121 (x - (-10))^2 + (y - (-15))^2 = 11^2 Compare the last equation to (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2 = a^2\] e\where (h,k) is the the centre and r is the radius

jimthompson5910 (jim_thompson5910):

Keep in mind that (x-h)^2+(y-k)^2=r^2 has a center of (h,k) and a radius of r

OpenStudy (anonymous):

so the radius is 121 i'm guessing...and the center must be 10.-15?

OpenStudy (anonymous):

10,-15

jimthompson5910 (jim_thompson5910):

the radius is NOT 121

jimthompson5910 (jim_thompson5910):

r^2 = 121 r = ???

OpenStudy (anonymous):

11 ^w^!

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

(h,k) = ???

OpenStudy (anonymous):

10,15?

jimthompson5910 (jim_thompson5910):

close, but your signs are off

OpenStudy (anonymous):

-10,-15

jimthompson5910 (jim_thompson5910):

if x - h = x + 10 then h must be -10

jimthompson5910 (jim_thompson5910):

yep, center is (-10, -15)

OpenStudy (anonymous):

yes! ^w^ thanks dude!

jimthompson5910 (jim_thompson5910):

np